find the distance between each pair of points
(-3,0) and (4,0)

Answer by checkley71(8403)

(-3,0) & (4,0)
X DISTANCE=4+3=7 ANSWER.
Y DISTANCE=0+0=0 ANSWER.
PROOF WOULD BE VERY EASY BY PLOTING THESE 2 POINTS ON GRAPH PAPER.